# Meta Curvature-Aware Minimization for Domain Generalization

### v1

@@ 3.2 Curvature Metric
Since SAM measures sharpness based solely on the loss value, it may fail to accurately capture the true sharpness in some situations [83, 89, 75]. As shown in Figure 2 (a), a smaller $f(\theta^{sam})$ does not always guarantee a flatter minima. Additionally, SAM consistently minimizes sharpness during training, without considering that the model should focus on minimizing the loss during early stages of training and only reduce sharpness near local minima. We propose a more rational training process that follows two principles: (1) minimize the loss when the model is far from convergence; and (2) reduce the sharpness when the model approaches a local minimum. Inspired by [18], we define a curvature metric that is independent of the loss value and adheres to these principles, providing a more accurate measure of sharpness and improving the model’s training process, as shown in Figure 2 (b). The curvature metric is formulated as:
$\mathcal{C}(f(\theta))=\frac{H(\theta)}{\|\nabla f(\theta)\|^{2}+1}.$ (2)
This metric plays a minor role when the model is far from convergence but becomes more significant as the model approaches a local minimum. We also explore the relationship between the curvature metric and the maximal eigenvalue of the Hessian, $\lambda_{max}(H(\theta))$, which is a proper measure of the sharpness of the local minima [28, 29] and is related to the generalization ability of the model [11]. Although $\lambda_{max}(H(\theta))$ is challenging to approximate and optimize directly [83, 78], we derive the following lemma showing that our curvature metric serves as a suitable surrogate for $\lambda_{max}(H(\theta))$.
Figure 2: Comparison between SAM and the proposed curvature metric. (a) Illustration of a sharp local minimum at $\theta_{1}$ (red) and a flat local minimum at $\theta_{2}$ (blue). While SAM prefers to optimize around $\theta_{2}^{sam}$, $\theta_{2}$ is inherently flatter than $\theta_{1}$. This highlights SAM’s limitation in accurately characterizing sharpness, as it is influenced by the loss value. (b) Illustration of the proposed curvature metric $\mathcal{C}$. Unlike SAM, $\mathcal{C}$ more effectively quantifies the deviation from a flat surface, providing a better measure of sharpness in the loss landscape. A smaller value of $\mathcal{C}$ indicates a flatter minimum. Best viewed in color.
Lemma 3.1. Suppose that $\|\mathcal{C}(f(\theta))\|$ is sub-multiplicative. Then, we have
$\lambda_{max}(H(\theta))=\|\mathcal{C}(f(\theta))\|\cdot\|\|\nabla f(\theta)\|^{2}+1\|.$ (3)
The proof can be found in Supplementary Material A.
@@ 3.3 Meta Curvature-Aware Minimization
As $\theta$ approaches a local minimum, the gradient $\nabla f(\theta)$ in Eq. (2) inevitably decreases. Consequently, we minimize the Hessian to reduce $\mathcal{C}$. Based on this analysis, we propose the MeCAM framework, which optimizes the following overall objective:
$\min_{\theta}f(\theta)+H(\theta),$ (4)
where minimizing $f(\theta)$ helps $\theta$ approach the local minima, and minimizing $H(\theta)$ focuses on reducing the curvature around the local minima.
@@
$H(\theta)=\frac{f(\theta+h)+f(\theta-h)-2f(\theta)}{h^{2}}+O,$ (5)
where $h$ is a small step-size, and $O$ denotes the remainder term. We estimate $H(\theta)$ by calculating finite differences at symmetric neighboring points around the point of interest. By incorporating information from both sides, this approach can provide an accurate approximation of $H(\theta)$. A key consideration in Eq. (5) is the choice of $h$. Common strategies for choosing an appropriate value of $h$ include balancing round-off errors, adjusting based on the scale of the problem, using typical small constants, and so on [55, 14]. To avoid additional computation, in this study, we empirically set
$h=\delta=$ $\rho\frac{\nabla f(\theta)}{\|\nabla f(\theta)\|+\epsilon},$ (6)
$\frac{1}{h^{2}}=\alpha,$
where $\alpha$ is a hyperparameter. The objective in Eq. (4) can then be rewritten as:
@@
where $\beta$ is another hyperparameter balancing the terms, with the condition of $\beta\leq\alpha$. The proof can be found in Supplementary Material A. Since we denote $f(\theta;\mathcal{D})$ as $f(\theta)$, we utilize $m(\theta-\delta)$ to represents $f(\theta-\delta;\tilde{\mathcal{D}})$ for clarity. $m(\theta-\delta)$ is designed to emulate the deployment environment to assess the model when transferred from the training set $\mathcal{D}$ to the virtual meta-test set $\tilde{\mathcal{D}}$, indicating the robustness of the current model across different data distributions. Finally, by combining both Eq. (7) and Eq. (10), the overall objective of MeCAM is:
@@ 4.2 Comparison with Existing DG Methods
We compared our MeCAM with 25 existing conventional DG methods [40, 85, 62, 4, 81, 33, 43, 64, 21, 26, 7, 77, 39, 70, 66, 52, 31, 67, 8, 10, 13] to evaluate its generalization ability. The results were presented in Table 2. Compared to the ERM algorithm [70], which can be treated as the baseline, MeCAM consistently demonstrates a substantial improvement across all benchmarks, with an average performance gain of $2.7\%$. MeCAM outperforms most methods on all datasets, with particularly notable improvements on PACS, VLCS, and DomainNet. Specifically, MeCAM achieves a $1.8\%$ improvement over MADG on PACS, a $1.1\%$ improvement over GMDG on VLCS, and a $0.4\%$ improvement over GMDG on DomainNet. Although MADG achieves the highest performance on OfficeHome and TerraIncognita, its performance on DomainNet is relatively weak. DomainNet is the largest dataset with over half a million images across 345 categories and collected from six distinct domains, posing a significant challenge. Our MeCAM surpasses other DG methods on this dataset, highlighting its robustness in handling such a large-scale and complex dataset. Overall, these results demonstrate that MeCAM improves the model generalization across a wide range of DG tasks.
@@ 4.3 Comparison with Optimizers and Sharpness-Based DG Methods
Table 4: Ablation study on various combinations of optimization objectives in our MeCAM. Average accuracy and standard error ($mean_{\pm std}$) of our MeCAM and other variants are calculated across three trials on PACS, VLCS, and OfficeHome. The best results are highlighted in bold.
$85.0_{\pm 0.5}$ $78.2_{\pm 0.5}$ $67.8_{\pm 0.2}$ $77.0$
$15.4_{\pm 0.3}$ $47.0_{\pm 0.3}$ $1.6_{\pm 0.1}$ $21.3$
$46.1_{\pm 5.2}$ $44.2_{\pm 2.9}$ $2.3_{\pm 0.2}$ $30.9$
$86.4_{\pm 1.3}$ $\textbf{80.3}_{\pm 0.5}$ $69.7_{\pm 0.2}$ $78.8$
$86.7_{\pm 0.5}$ $79.7_{\pm 0.3}$ $69.5_{\pm 0.5}$ $78.6$
$87.4_{\pm 0.1}$ $79.4_{\pm 0.1}$ $69.8_{\pm 0.2}$ $78.9$
$\textbf{88.3}_{\pm 0.3}$ $\textbf{80.3}_{\pm 0.3}$ $\textbf{70.4}_{\pm 0.4}$ 79.7
4.4 Ablation Study
To better understand the contributions of each component in MeCAM, we conducted an ablation study on the three terms of MeCAM’s optimization objective (see Eq. (11)). The results were displayed in Table 4. Please note that when only $\alpha(f(\theta+\delta)-f(\theta))$ or $\beta(m(\theta-\delta)-f(\theta))$ is used as the optimization objective, we set $\alpha$ or $\beta$ to 1. The results highlight the following insights: (1) using only $f(\theta)$ degenerates MeCAM to ERM, which does not effectively reduce sharpness or curvature; (2) relying solely on the surrogate gap of SAM or meta-learning results in significant performance degradation, particularly on the OfficeHome dataset, indicating that minimizing sharpness alone is insufficient; (3) combining the surrogate gaps of SAM or meta-learning with $f(\theta)$ alleviates this degradation and improves generalization performance; and (4) the best performance is achieved when all three terms are jointly used in the optimization objective, suggesting the effectiveness of MeCAM. Furthermore, it is worth noting that combining the surrogate gap of SAM with meta-learning not only resolves the degradation observed when using either of them individually, but also outperforms the use of $f(\theta)$ alone. We argue that the optimization objectives of these two terms are complementary, and their combination effectively mitigates limitations of each other.
@@ 4.5 Evaluation on Various Training Iterations
4.6 Curvature Comparison on PACS
In this experiment, we evaluated the curvature metric $\mathcal{C}$ around the local minima obtained by our MeCAM and other competing methods. Based on the above approximation, we utilized various values of $\rho$ to compute $\mathcal{C}$ on the PACS dataset by $\mathcal{C}(f(\theta))=\alpha\frac{f(\theta+\delta)+f(\theta-\delta)-2f(\theta)}{\|\nabla f(\theta)\|^{2}+1}$, where $\alpha$ is omitted for convenient comparison. The results were listed in Table 5, with smaller curvature indicating superior algorithm. It can be found that our MeCAM consistently outperforms other competing methods across most values of $\rho$ and improves $\mathcal{C}$ significantly as $\rho$ increases, demonstrating the robustness of our MeCAM against large perturbations.
@@ 4.6 Curvature Comparison on PACS
[…]
@@ Appendix A Proofs
A.1 Proof of Lemma 3.1
Proof. Since $\|\mathcal{C}(f(\theta))\|$ is sub-multiplicative, we have:
$\|\mathcal{C}(f(\theta))\|$ $\leq\frac{\|H(\theta)\|}{\|\|\nabla f(\theta)\|^{2}+1\|}$ (12)
$=\frac{\lambda_{max}(H(\theta))}{\|\|\nabla f(\theta)\|^{2}+1\|}$
$\lambda_{max}(H(\theta))$ $\geq\|\mathcal{C}(f(\theta))\|\cdot\|\|\nabla f(\theta)\|^{2}+1\|.$
In addition,
$\|\mathcal{C}(f(\theta))\|\cdot\|\|\nabla f(\theta)\|^{2}+1\|$ $\geq\|\mathcal{C}(f(\theta))(\|\nabla f(\theta)\|^{2}+1)\|$ (13)
$=\|H(\theta)\|$
$=\lambda_{max}(H(\theta)).$
Combine Eq. (12) and Eq. (13), we have:
$\|\mathcal{C}(f(\theta))\|\cdot\|\|\nabla f(\theta)\|^{2}+1\|=\lambda_{max}(H(\theta)).$ (14)
A.2 Proof of Proposition 3.3
@@
A.3.1 Convergence W.R.T. Function $f(\theta)$
@@
A.3.2 Convergence W.R.T. Function $f(\theta^{sam})$
@@
A.3.3 Convergence W.R.T. Function $f(\theta^{meta})$
@@
Similar to the proof in A.3.2, we take the expectation conditioned on the observations up to step $t$ for the both sides and obtain:
@@
A.3.4 Convergence W.R.T. Overall Function $F(\theta)$
Note that we have proved convergence for $f(\theta)$, $f(\theta^{sam})$, and $f(\theta^{meta})$ in A.3.1, A.3.2, and A.3.3, respectively. By definition of $F(\theta)$, we have:
@@
A.4 Proof of Proposition 3.5
@@
Table I: The optimal hyperparameter configurations for MeCAM on each dataset. The ‘lr’, ‘wd’, ‘dr’, $\rho$, $\alpha$, and $\beta$ denote the learning rate, the weight decay, the dropout rate, the perturbation radius, the weight for the surrogate gap of SAM, and the weight for the surrogate gap of meta-learning, respectively.
  Dataset lr wd dr $\rho$ $\alpha$ $\beta$
PACS $3e^{-5}$ $1e^{-4}$ $0.5$ $0.1$ $0.1$ $0.1$
VLCS $1e^{-6}$ $1e^{-4}$ $0.5$ $0.1$ $0.2$ $0.1$
OfficeHome $1e^{-5}$ $1e^{-4}$ $0.5$ $0.2$ $0.2$ $0.1$
TerraIncognita $1e^{-5}$ $1e^{-4}$ $0.5$ $0.01$ $0.05$ $0.05$
DomainNet $3e^{-5}$ $1e^{-6}$ $0.5$ $0.1$ $0.2$ $0.1$
@@ Appendix C Full Results
In this section, we provide the detailed results of Table 2 and Table 3 in our manuscript. Specifically, we display the full results of our MeCAM and existing DG methods [40, 85, 62, 4, 81, 33, 56, 43, 21, 26, 7, 77, 39, 70, 52, 67, 60, 13, 69] on PACS, VLCS, OfficeHome, TerraIncognita, and DomainNet datasets in Table II, Table IV, Table VI, Table VIII, and Table X, respectively. We also display the full results of our MeCAM, optimizers [32, 47, 53, 80, 88, 79], and existing sharpness-based DG methods [20, 83, 75, 82, 76, 42] on PACS, VLCS, OfficeHome, TerraIncognita, and DomainNet datasets in Table III, Table V, Table VII, Table IX, and Table XI, respectively. The average accuracy and standard deviation, calculated across three trials with random seeds set to 0, 1, and 2, are reported for each method.
@@
MMD [40] $60.4_{\pm 0:2}$ $53.3_{\pm 0:3}$ $74.3_{\pm 0:1}$ $77.4_{\pm 0:6}$ $66.4$
Mixstyle [85] $51.1_{\pm 0:3}$ $53.2_{\pm 0:4}$ $68.2_{\pm 0:7}$ $69.2_{\pm 0:6}$ $60.4$
GroupDRO [62] $60.4_{\pm 0:7}$ $52.7_{\pm 1:0}$ $75.0_{\pm 0:7}$ $76.0_{\pm 0:7}$ $66.0$
IRM [4] $58.9_{\pm 2:3}$ $52.2_{\pm 1:6}$ $72.1_{\pm 2:9}$ $74.0_{\pm 2:5}$ $64.3$
ARM [81] $58.9_{\pm 0:8}$ $51.0_{\pm 0:5}$ $74.1_{\pm 0:1}$ $75.2_{\pm 0:3}$ $64.8$
VREx [33] $60.7_{\pm 0:9}$ $53.0_{\pm 0:9}$ $75.3_{\pm 0:1}$ $76.6_{\pm 0:5}$ $66.4$
@@
CDANN [43] $61.5_{\pm 1:4}$ $50.4_{\pm 2:4}$ $74.4_{\pm 0:9}$ $76.6_{\pm 0:8}$ $65.7$
DANN [21] $59.9_{\pm 1:3}$ $53.0_{\pm 0:3}$ $73.6_{\pm 0:7}$ $76.9_{\pm 0:5}$ $65.9$
RSC [26] $60.7_{\pm 1:4}$ $51.4_{\pm 0:3}$ $74.8_{\pm 1:1}$ $75.1_{\pm 1:3}$ $65.5$
MTL [7] $61.5_{\pm 0:7}$ $52.4_{\pm 0:6}$ $74.9_{\pm 0:4}$ $76.8_{\pm 0:4}$ $66.4$
@@
MMD [40] $41.9_{\pm 3:0}$ $34.8_{\pm 1:0}$ $57.0_{\pm 1:9}$ $35.2_{\pm 1:8}$ $42.2$
Mixstyle [85] $54.3_{\pm 1:1}$ $34.1_{\pm 1:1}$ $55.9_{\pm 1:1}$ $31.7_{\pm 2:1}$ $44.0$
GroupDRO [62] $41.2_{\pm 0:7}$ $38.6_{\pm 2:1}$ $56.7_{\pm 0:9}$ $36.4_{\pm 2:1}$ $43.2$
IRM [4] $54.6_{\pm 1:3}$ $39.8_{\pm 1:9}$ $56.2_{\pm 1:8}$ $39.6_{\pm 0:8}$ $47.6$
ARM [81] $49.3_{\pm 0:7}$ $38.3_{\pm 2:4}$ $55.8_{\pm 0:8}$ $38.7_{\pm 1:3}$ $45.5$
VREx [33] $48.2_{\pm 4:3}$ $41.7_{\pm 1:3}$ $56.8_{\pm 0:8}$ $38.7_{\pm 3:1}$ $46.4$
@@
CDANN [43] $47.0_{\pm 1:9}$ $41.3_{\pm 4:8}$ $54.9_{\pm 1:7}$ $39.8_{\pm 2:3}$ $45.8$
DANN [21] $51.1_{\pm 3:5}$ $40.6_{\pm 0:6}$ $57.4_{\pm 0:5}$ $37.7_{\pm 1:8}$ $46.7$
RSC [26] $50.2_{\pm 2:2}$ $39.2_{\pm 1:4}$ $56.3_{\pm 1:4}$ $40.8_{\pm 0:6}$ $46.6$
MTL [7] $49.3_{\pm 1:2}$ $39.6_{\pm 6:3}$ $55.6_{\pm 1:1}$ $37.8_{\pm 0:8}$ $45.6$

### v2

@@ 3.2 Curvature Metric
Since SAM measures sharpness based solely on the loss value, it may fail to accurately capture the true sharpness in some situations [82, 88, 74]. As shown in Figure 2 (a), a smaller $f(\theta^{sam})$ does not always guarantee a flatter minima. Additionally, SAM consistently minimizes sharpness during training, without considering that the model should focus on minimizing the loss during early stages of training and only reduce sharpness near local minima. We propose a more rational training process that follows two principles: (1) minimize the loss when the model is far from convergence; and (2) reduce the sharpness when the model approaches a local minimum. Inspired by [18], we define a curvature metric that is independent of the loss value and adheres to these principles, providing a more accurate measure of sharpness and improving the model’s training process. The curvature metric is formulated as:
$\mathcal{C}(f(\theta))=\frac{\|H(\theta)\|}{\|\nabla f(\theta)\|^{2}+1}.$ (2)
This metric plays a minor role when the model has not converged but becomes more significant as the model approaches a local minimum, as shown in Figure 2 (b). We also explore the relationship between this metric and the maximal eigenvalue of the Hessian, $\lambda_{max}(H(\theta))$, which is a proper measure of the sharpness of the local minima [28, 29] and is related to the generalization ability of the model [11]. Although $\lambda_{max}(H(\theta))$ is challenging to approximate and optimize directly [82, 77], we derive the following lemma showing that our curvature metric can be served as a suitable surrogate for it.
Lemma 3.1. For a local minima, assuming that $H(\theta)$ is positive semi-definite [88], we have
$\lambda_{max}(H(\theta))=\mathcal{C}(f(\theta)).$ (3)
This indicates that $\lambda_{max}(H(\theta))$ can be directly controlled by our curvature metric.
Figure 2: Comparison between SAM and the proposed curvature metric. (a) Illustration of a sharp local minimum at $\theta_{1}$ (red) and a flat local minimum at $\theta_{2}$ (blue), where $S(\theta)$ denotes the sharpness of model with parameter $\theta$. While SAM prefers to optimize around $\theta_{2}^{sam}$, $\theta_{2}$ is inherently flatter than $\theta_{1}$. This highlights SAM’s limitation in accurately characterizing sharpness, as it is influenced by the loss value. (b) Illustration of the proposed curvature metric $\mathcal{C}$. Unlike SAM, $\mathcal{C}$ more effectively quantifies the deviation from a flat surface, providing a better measure of sharpness in the loss landscape. A smaller value of $\mathcal{C}$ indicates a flatter minimum. Best viewed in color.
@@ 3.3 Meta Curvature-Aware Minimization
As $\theta$ approaches a local minimum, the gradient $\nabla f(\theta)$ in Eq. (2) inevitably decreases. Consequently, we minimize the Hessian to reduce the numerator approximatively. Based on this analysis, we propose the MeCAM framework, which optimizes the following overall objective:
$\min_{\theta}f(\theta),H(\theta).$ (4)
Here, minimizing $f(\theta)$ helps $\theta$ approach the local minima, and minimizing $H(\theta)$ focuses on reducing the curvature around the local minima.
@@
$H(\theta)\approx\frac{f(\theta+h)+f(\theta-h)-2f(\theta)}{h^{2}},$ (5)
where $h$ is a small step-size, and the remainder term is omitted. We estimate $H(\theta)$ by calculating finite differences at symmetric neighboring points around the point of interest. By incorporating information from both sides, this approach can provide an accurate approximation of $H(\theta)$. A key consideration in Eq. (5) is the choice of $h$. Common strategies for choosing an appropriate value of $h$ include balancing round-off errors, adjusting based on the scale of the problem, using typical small constants, and so on [55, 14]. To avoid additional computation, in this study, we empirically set
$h=\delta=$ $\rho\frac{\nabla f(\theta)}{\|\nabla f(\theta)\|+\epsilon}$ (6)
and utilize a hyperparameter $\alpha$ to replace $\frac{1}{h^{2}}$ to better control the contribution of this term during training. The objective in Eq. (4) can then be rewritten as:
@@
When the model has not converged, assuming $f$ is monotonic and smooth within $[\theta+\delta,\theta-\delta]$ and $\delta$ is small, the sum of the second and third items in Eq. 7 tends to 00. Conversely, as the model approaches a local minimum, this sum remains nonnegative and is positively correlated with the model’s sharpness.
@@
where $\beta$ is another hyperparameter balancing the terms, with the condition of $\beta\leq\alpha$. The proof can be found in Supplementary Material A. Since we denote $f(\theta;\mathcal{D})$ as $f(\theta)$, we utilize $m(\theta-\delta)$ to represents $f(\theta-\delta;\tilde{\mathcal{D}})$ for clarity. $m(\theta-\delta)$ is designed to emulate the deployment environment to assess the model when transferred from the training set $\mathcal{D}$ to the virtual meta-test set $\tilde{\mathcal{D}}$, indicating the robustness of the current model across different data distributions. Finally, by combining both Eq. (7) and Eq. (10), the overall objective of MeCAM can be defined as:
@@ 4.2 Comparison with Existing DG Methods
We compared our MeCAM with 22 existing conventional DG methods [40, 84, 62, 4, 80, 33, 43, 64, 21, 7, 76, 39, 69, 66, 52, 31, 8, 60, 10, 34, 13] to evaluate its generalization ability. The results were presented in Table 2. Compared to the ERM algorithm [69], which can be treated as the baseline, MeCAM consistently demonstrates a substantial improvement across all benchmarks, with an average performance gain of $2.7\%$. MeCAM outperforms most methods on all datasets, with particularly notable improvements on PACS, VLCS, and DomainNet. Specifically, MeCAM achieves a $1.8\%$ improvement over MADG on PACS, a $1.3\%$ improvement over MIRO on VLCS, and a $0.7\%$ improvement over MIRO on DomainNet. Although MADG achieves the highest performance on OfficeHome and TerraIncognita, its performance on DomainNet is relatively weak. DomainNet is the largest dataset with over half a million images across 345 categories and collected from six distinct domains, posing a significant challenge. Our MeCAM surpasses other DG methods on this dataset, highlighting its robustness in handling such a large-scale and complex dataset. Overall, these results demonstrate that MeCAM improves the model generalization across a wide range of DG tasks.
@@ 4.3 Comparison with Optimizers and Sharpness-Based DG Methods
[…]
  Algorithm PACS VLCS OfficeHome TerraInc DomainNet Average
Adam + SWAD $86.8$ $79.1$ $70.1$ $46.5$ $44.1$ $65.3$
AdamW + SWAD $87.0$ $78.5$ $70.8$ $46.9$ $45.0$ $65.6$
SGD + SWAD $85.2$ $79.1$ $71.0$ $46.7$ $42.8$ $65.0$
FAD + SWAD $88.5$ 79.8 71.8 $47.5$ $45.0$ $66.5$
MeCAM (Ours) + SWAD 88.9 79.8 $70.8$ 48.9 45.9 66.9
Adam + CORAL $86.0$ $78.9$ $68.7$ $43.7$ $44.5$ $64.5$
AdamW + CORAL $86.4$ $79.5$ $69.8$ $45.0$ $44.9$ $65.1$
SGD + CORAL $85.6$ $78.2$ $69.5$ $45.8$ $44.6$ $64.7$
FAD + CORAL 88.5 $78.9$ $70.8$ $46.1$ $44.9$ $65.9$
MeCAM (Ours) + CORAL $88.0$ 79.7 71.0 49.9 45.0 66.7
Adam + RSC $84.5$ $77.9$ $65.7$ $44.5$ $42.8$ $63.1$
AdamW + RSC $83.4$ $77.5$ $66.3$ $45.1$ $42.4$ $62.9$
SGD + RSC $82.6$ $78.1$ $67.0$ $43.9$ $43.5$ $63.0$
FAD + RSC $86.9$ $77.6$ $68.6$ $46.2$ $44.1$ $64.7$
MeCAM (Ours) + RSC 87.2 79.8 70.2 47.3 44.7 65.8
4.4 Extensibility Analysis
To evaluate the extensibility of our MeCAM, we integrated it with three representative DG methods (i.e., SWAD [9], CORAL [67], and RSC [26]), and repeated experiments on five DG datasets. The optimal hyperparameter configurations listed in Supplementary Material B were directly applied to MeCAM during its integration with other DG methods. Therefore, no additional hyperparameter search was performed. The results were presented in Table 4. It shows that the combinations involved our MeCAM consistently outperforms others across most scenarios and also achieve the best overall performance, demonstrating the superior extensibility of our MeCAM.
@@ 4.5 Evaluation on Various Training Iterations
4.6 Curvature Comparison on PACS
In this experiment, we evaluated the curvature metric $\mathcal{C}$ around the local minima obtained by our MeCAM and other competing methods. Based on the above approximation, we utilized various values of $\rho$ to compute $\mathcal{C}$ on the PACS dataset by $\mathcal{C}(f(\theta))=\alpha\frac{|f(\theta+\delta)+f(\theta-\delta)-2f(\theta)|}{\|\nabla f(\theta)\|^{2}+1}$, where $\alpha$ is omitted for convenient comparison. The results were listed in Table 5, with smaller curvature indicating superior algorithm. It can be found that our MeCAM consistently outperforms other competing methods across all values of $\rho$ and improves $\mathcal{C}$ significantly as $\rho$ increases, demonstrating the robustness of our MeCAM against large perturbations.
@@ Appendix A Proofs
A.1 Proof of Proposition 3.3
@@
A.2.1 Convergence W.R.T. Function $f(\theta)$
@@
A.2.2 Convergence W.R.T. Function $f(\theta^{sam})$
@@
A.2.3 Convergence W.R.T. Function $f(\theta^{meta})$
@@
Similar to the proof in A.2.2, we take the expectation conditioned on the observations up to step $t$ for the both sides and obtain:
@@
A.2.4 Convergence W.R.T. Overall Function $F(\theta)$
Note that we have proved convergence for $f(\theta)$, $f(\theta^{sam})$, and $f(\theta^{meta})$ in A.2.1, A.2.2, and A.2.3, respectively. By definition of $F(\theta)$, we have:
@@
Table I: The optimal hyperparameter configurations for MeCAM on each dataset. The $\rho$, $\alpha$, and $\beta$ denote the perturbation radius, the weight for the surrogate gap of SAM, and the weight for the surrogate gap of meta-learning, respectively.
  Dataset learning rate weight decay dropout rate $\rho$ $\alpha$ $\beta$
PACS $3e^{-5}$ $1e^{-4}$ $0.5$ $0.1$ $0.1$ $0.1$
VLCS $1e^{-6}$ $1e^{-4}$ $0.5$ $0.1$ $0.2$ $0.1$
OfficeHome $1e^{-5}$ $1e^{-4}$ $0.5$ $0.2$ $0.2$ $0.1$
TerraIncognita $1e^{-5}$ $1e^{-4}$ $0.5$ $0.01$ $0.05$ $0.05$
DomainNet $3e^{-5}$ $1e^{-6}$ $0.5$ $0.1$ $0.2$ $0.1$
A.3 Proof of Proposition 3.5
@@ Appendix C Full Results
In this section, we provide the detailed results of Table 2 and Table 3 shown in our manuscript. Specifically, we display the full results of our MeCAM and other existing DG methods [40, 84, 62, 4, 80, 33, 56, 43, 21, 7, 76, 39, 69, 52, 60, 13] on PACS, VLCS, OfficeHome, TerraIncognita, and DomainNet datasets in Table II, Table IV, Table VI, Table VIII, and Table X, respectively. We also display the full results of our MeCAM, optimizers [32, 47, 53, 79, 87, 78], and existing sharpness-based DG methods [20, 82, 74, 81, 75, 42] on PACS, VLCS, OfficeHome, TerraIncognita, and DomainNet datasets in Table III, Table V, Table VII, Table IX, and Table XI, respectively. For the results of each method, we report the average accuracy and standard deviation computed over three trials using random seeds of 0, 1, and 2.
@@
MMD [40] $60.4_{\pm 0.2}$ $53.3_{\pm 0.3}$ $74.3_{\pm 0.1}$ $77.4_{\pm 0.6}$ $66.4$
Mixstyle [84] $51.1_{\pm 0.3}$ $53.2_{\pm 0.4}$ $68.2_{\pm 0.7}$ $69.2_{\pm 0.6}$ $60.4$
GroupDRO [62] $60.4_{\pm 0.7}$ $52.7_{\pm 1.0}$ $75.0_{\pm 0.7}$ $76.0_{\pm 0.7}$ $66.0$
IRM [4] $58.9_{\pm 2.3}$ $52.2_{\pm 1.6}$ $72.1_{\pm 2.9}$ $74.0_{\pm 2.5}$ $64.3$
ARM [80] $58.9_{\pm 0.8}$ $51.0_{\pm 0.5}$ $74.1_{\pm 0.1}$ $75.2_{\pm 0.3}$ $64.8$
VREx [33] $60.7_{\pm 0.9}$ $53.0_{\pm 0.9}$ $75.3_{\pm 0.1}$ $76.6_{\pm 0.5}$ $66.4$
@@
CDANN [43] $61.5_{\pm 1.4}$ $50.4_{\pm 2.4}$ $74.4_{\pm 0.9}$ $76.6_{\pm 0.8}$ $65.7$
SAND-mask [64] $59.9_{\pm 0.7}$ $53.6_{\pm 0.8}$ $74.3_{\pm 0.4}$ $75.8_{\pm 0.5}$ $65.9$
DANN [21] $59.9_{\pm 1.3}$ $53.0_{\pm 0.3}$ $73.6_{\pm 0.7}$ $76.9_{\pm 0.5}$ $65.9$
MTL [7] $61.5_{\pm 0.7}$ $52.4_{\pm 0.6}$ $74.9_{\pm 0.4}$ $76.8_{\pm 0.4}$ $66.4$
@@
MMD [40] $41.9_{\pm 3.0}$ $34.8_{\pm 1.0}$ $57.0_{\pm 1.9}$ $35.2_{\pm 1.8}$ $42.2$
Mixstyle [84] $54.3_{\pm 1.1}$ $34.1_{\pm 1.1}$ $55.9_{\pm 1.1}$ $31.7_{\pm 2.1}$ $44.0$
GroupDRO [62] $41.2_{\pm 0.7}$ $38.6_{\pm 2.1}$ $56.7_{\pm 0.9}$ $36.4_{\pm 2.1}$ $43.2$
IRM [4] $54.6_{\pm 1.3}$ $39.8_{\pm 1.9}$ $56.2_{\pm 1.8}$ $39.6_{\pm 0.8}$ $47.6$
ARM [80] $49.3_{\pm 0.7}$ $38.3_{\pm 2.4}$ $55.8_{\pm 0.8}$ $38.7_{\pm 1.3}$ $45.5$
VREx [33] $48.2_{\pm 4.3}$ $41.7_{\pm 1.3}$ $56.8_{\pm 0.8}$ $38.7_{\pm 3.1}$ $46.4$
@@
CDANN [43] $47.0_{\pm 1.9}$ $41.3_{\pm 4.8}$ $54.9_{\pm 1.7}$ $39.8_{\pm 2.3}$ $45.8$
SAND-mask [64] $56.2_{\pm 1.8}$ $46.3_{\pm 0.3}$ $55.8_{\pm 0.4}$ $42.6_{\pm 1.2}$ $50.2$
DANN [21] $51.1_{\pm 3.5}$ $40.6_{\pm 0.6}$ $57.4_{\pm 0.5}$ $37.7_{\pm 1.8}$ $46.7$
MTL [7] $49.3_{\pm 1.2}$ $39.6_{\pm 6.3}$ $55.6_{\pm 1.1}$ $37.8_{\pm 0.8}$ $45.6$